OpenStudy (anonymous):

Can someone check my work please? Jared works for Bloom's Landscaping and will be planting a flower garden with rows of flowers. According to the plan, he will plant 6 plants in Row 1, 8 plants in Row 2, 10 plants in Row 3, and so on. A) How many plants will be planted in Row 10? Show your work. a) Rows 1/2/ 3 / 4 /5/ 6 / 7/ 8/ 9 / 10/11/12 Plants ->/6/8/10/12/14/16/18/20/22/24/26/28 __________________________________________________ B) If the landscaper wants Jared to plant 12 rows of plants, how many plants will be needed? b) 204 plants will be needed. (to be continued)

OpenStudy (anonymous):

C) If the landscaper purchased 300 plants, how many rows would JAred be able to make? How many plants, if any, would be left over? C) I dont know how to do this one. _________________________________________________________________________________________ D) If Jared started Row 1 with twice as many plants, (12 plants) and increased the number by 2 in each row, will Jared need twice as many plants as were needed in the original plan if 12 rows are planted? d) i dont know how to do this one either.

OpenStudy (phi):

how did you do part A?

OpenStudy (anonymous):

I just went by two's @phi

OpenStudy (phi):

do you know the formula for adding up N numbers: 0+1+2+3+...+N= N*(N+1)/2 if you go up by 2's we can tweak that formula

OpenStudy (phi):

example: 0+1+2 here N=2 (last number) so the formula says the sum should be N*(N+1)/2 = 2*3/2= 3 and 1+2 = 3 it works. It will also work for all N

OpenStudy (phi):

if we go up by 2: 0+2+4+6 that is like: 2(0+1+2+3) = 2*6= 12 so 2* sum of numbers from 0 to 3= 2* 6=12 is the fast way.

OpenStudy (anonymous):

So for B it would be 12(6+28) divide it by 2 then 12(17) = 204?

OpenStudy (phi):

if you have 12 rows: 6+0 + (6+2) + (6+4) + ... (6+22) we can write that sum as 6+6+...+6 + 0+2+4+...+22 we have 12 6's so that is 6*12 + 2(0+1+2+...+11) simplify that

OpenStudy (phi):

yes for B

OpenStudy (anonymous):

I think for D we just multiply 204 with two right?

OpenStudy (anonymous):

72 + 28 = 100

OpenStudy (phi):

I think there is a formula to find the last term in a sequence. then do the sum formula. In other words, I would figure it out rather than try to intuit.

OpenStudy (phi):

but it is clear if we start with 6 extra plants, we will have 6*12= 72 extra plants total added to the answer for part B

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you lol

OpenStudy (phi):

Did you get part C ?

OpenStudy (anonymous):

I think its 24 rows of plants would be needed and 52 left over?

OpenStudy (anonymous):

@phi will you help me with a different question?

OpenStudy (phi):

in a minute

OpenStudy (phi):

the simplified formula for the sum of N rows (for your problem only) seems to be (N+5)*N so for 1 row: (1+5)*1= 6 works for 2 rows : (2+5)*2 = 14 = 6+8 works for 12 rows: (12+5)*12= 17*12= 204 so for part C: (N+5)*N= 300 N^2 +5N-300 = 0 (N+20)(N-15)= 0 N= -20 can't use this N= 15 this one works

OpenStudy (phi):

in other words, exactly 15 rows, with no plants left over.